OpenStudy (anonymous):

for a circle of radius 2 feet, find the arc length s subtended by a central angle of 54 degrees

jimthompson5910 (jim_thompson5910):

What's the circumference of the full circle?

OpenStudy (anonymous):

Use the formula: [[(54º)*pi]/180º]*2ft

OpenStudy (anonymous):

54 degree = (54 * pi)/180 radian = 0.942 radian Arch Length = Radius * central angle = 2* 0.942 = 1.884 feet.